OpenStudy (anonymous):

The table shows the number of minutes customers waited in line at a bank during the lunch hour, and the corresponding frequencies. Minutes spent waiting Frequency 1 3 2 10 3 2 4 1 5 4 6 or more 7 Find the probability of each event. (Round your answers to the nearest percent.) (a) A customer waited 4 minutes. % (b) A customer waited 3 or 4 minutes. % (c) A customer waited less than 4 minutes. % (d) A customer waited 5 or more minutes. %

OpenStudy (anonymous):

P(A) = The Number Of Ways Event A Can Occur ----------------------------------- The total number Of Possible Outcomes

OpenStudy (anonymous):

For parts a,b,c and d your denominator will be the same. What is your denominator?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

No your denominator for each part is just the total number of customer, so you have to add up 3+10+2+1+4+7

OpenStudy (anonymous):

27

OpenStudy (anonymous):

right. now lets look at part a. how many customers waited 4 minutes? That 's your numerator for part a.

OpenStudy (anonymous):

so 4/27= 0.14 or 14%?

OpenStudy (anonymous):

no. only 1 customer waited 4 minutes 1/27 = 0.04 or 4%

OpenStudy (anonymous):

1 customer out of 27 customers waited 4 minutes. make sense?

OpenStudy (anonymous):

yes, this makes sense.

OpenStudy (anonymous):

how many customers waited 3 or 4 minutes?

OpenStudy (anonymous):

3, so 3/27= 11% correct?

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

So, to determine how many students waited less than 4 minutes, do I use the numbers for 4 minutes and down or start at 3?

OpenStudy (anonymous):

Im confused with that one

OpenStudy (anonymous):

just 3 + 10 + 2

OpenStudy (anonymous):

so 15/27?

OpenStudy (anonymous):

15/27=55% correct? or is it rounded to 56%?

OpenStudy (anonymous):

round up since the next digit after the 2nd 5 is a 5 0.55555555 rounds to 0.56 or 56%

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

for 5 or more minutes use 7+4 .correct?

OpenStudy (anonymous):

and that would be 11/27 = 41%